OpenStudy (anonymous):

Two hard steel balls collide in a perfectly elastic collision. The first ball has a mass of 1kg and is traveling at 5m/s east. The second ball has a mass of 2kg and is initially traveling at 7 m/s west along the same path so they collide head on. After the collision the second ball is traveling at 1m/s east. What is the velocity of the 1st ball after the collision? **How do I solve and show the work for this? :/ Thanks!!

OpenStudy (michele_laino):

here we have to apply the law of conservation of total momentum and the law of conservation of kinetic energy, since the collision is elastic

OpenStudy (michele_laino):

the situation, before collision is: 

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so, total momentum is: \[\Large {m_1}{v_1} - {m_2}{v_2} = 1 \times 5 - 2 \times 7 = ...\]

OpenStudy (anonymous):

ok! so we get 5-14? = -9 ?

OpenStudy (michele_laino):

minus sign, means that the vector of the resultant momentum is oriented towards west

OpenStudy (anonymous):

ohh okay! so -9, meaning 9 m/s to the west is the velocity of the 1st ball after collision?

OpenStudy (michele_laino):

no, -9 is the magnitude and sign of the resultant momentum

OpenStudy (anonymous):

ohh okay! so it's like this? 5-14=-9 =resultant momentum ?

OpenStudy (michele_laino):

yes! now total kinetic energy is: \[\large KE = \frac{1}{2}{m_1}v_1^2 + \frac{1}{2}{m_2}v_2^2 = \frac{1}{2} \times 1 \times 25 + \frac{1}{2} \times 2 \times 49 = ...\]

OpenStudy (anonymous):

ok! so we get this? 61.5 ?

OpenStudy (michele_laino):

ok! correct!

OpenStudy (michele_laino):

now let's consider the situation after the collision

OpenStudy (anonymous):

ohhh so 61.5 is just the KE right? so we solve for velocity now? :/

OpenStudy (michele_laino):



OpenStudy (michele_laino):

no, we have to apply the conservation of total momentum first

OpenStudy (anonymous):

oh ok! how do we do that?

OpenStudy (michele_laino):

I call with u the speed of the first ball after collision

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so we can write: conservation of total momentum: \[\Large \begin{gathered} {m_1}u + {m_2}v{'_2} = - 9 \hfill \\ 1 \times u + 7 \times 1 = - 9 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

what is u?

OpenStudy (anonymous):

-16?

OpenStudy (michele_laino):

in my reasoning i have considered positive velocity if the velocity is towards east, and negative velocity if that velocity is towards west

OpenStudy (anonymous):

oh! so u=-16?

OpenStudy (anonymous):

16 m/s to the west?

OpenStudy (michele_laino):

correct! that means the first ball, after the collision is goinf towards west

OpenStudy (michele_laino):

going*

OpenStudy (anonymous):

ahh so that is our solution? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thank you!:D

OpenStudy (michele_laino):

:)